G=GMm/r² where f is magnitude of the force m and m are the sizes of each mass r is the distance between the masses and G is a constant if both M and r are halved keeping all other values constant what is the ratio of the new force F1 to the original force f0 (F1/f0


1
Expert's answer
2022-03-27T18:41:12-0400

Answer

Gravitation force is given

"f_0=\frac{GMm}{r^2}"

Now M and r is halved

So force "F_1=\frac{GMm}{2r^2/4}"

So ratio


"\frac{F_1}{f_0}=2"